OpenStudy (kl0723):

Set up the integral for the length of the curve:

OpenStudy (kl0723):

\[x=y^{\frac{ 1 }{ 7 }}, 0 \le y \le 2\]

OpenStudy (anonymous):

do u know the formul

OpenStudy (anonymous):

I know how to do the problem I just want to make sure that I have the correct answer.

OpenStudy (kl0723):

using \[\frac{ dx }{ dy } \] as in \[\int\limits \sqrt{1+(x^{'})^2} dy\]

OpenStudy (anonymous):

????? I'm lost

OpenStudy (kl0723):

well... tehre is no need to solve the integral... just to set up :)

OpenStudy (anonymous):

Okay thank you.

OpenStudy (aum):

\[ \int_0^2\sqrt{1+\left(\frac{dx}{dy}\right)^2}dy \]

OpenStudy (kl0723):

exactly... @aum pretty much derive to find x' and then plug in with the actual value for x

OpenStudy (kl0723):

x' = 1/7y^(-6/7)

OpenStudy (aum):

\[ x = y^{1/7} \\ \frac{dx}{dy} = \frac 17y^{-6/7} \]

OpenStudy (aum):

\[ \frac 17\int_0^2\sqrt{49+y^{-12/7}}dy \]

OpenStudy (kl0723):